If, in a monopoly market, the demand function for a product is p = 140 − 0.10x and the revenue function is R = px, where x is the number of units sold and p is the price per unit, what price will maximize revenue?
Answer #1

The revenue here is given as:

R = px = (140 - 0.1x)*x = 140x - 0.1x2

Differentiating it with respect to x, we get here:

dR d = 140 -0.2

Differentiating it again with respect to x, we get here:

dR d2 = -0.2 <0

Therefore we would get maximum value of R for dR / dx = 0

Therefore, we have here:

140 - 0.2x = 0

x = 140/0.2 = 700

Therefore the price that will maximize the revenue is computed as:

P = 140 - 0.1x = 140 - 0.1*700 = 70

Therefore 70 is the required price here.